Find the exact value of sin 75 degrees.
Make sure to show work.

Answer by Boreal(15235)

sin (a+b)=sinacosb+cosa*sinb
sin (45+30)=sin(75)
sin 45= sqrt(2)/2=cos 45
sin 3=1/2
cos30=sqrt(3)/2
sin 75=sqrt(2)/2 * sqrt(3)/2+ sqrt(2)/2*(1/2)=(1/4)*sqrt(6)+(1/4)sqrt(2)
=(1/4)(sqrt(6)+sqrt(2))